OpenStudy (anonymous):

Find the first four terms of the multiplication pattern given by the formula. 8 • 3n – 1

OpenStudy (loser66):

if n=1, then 3n =?

OpenStudy (anonymous):

3 ?

OpenStudy (loser66):

and then 3-1=?

OpenStudy (anonymous):

2

OpenStudy (loser66):

conclusion n=1, 3n-1 = 2 (the first term) do the same with n =2 3n-1 = ...(the second term) do the same with n =3 3n-1 =...(the third term) do the same with n =4 3n -1 =....(the fourth term) then arrange them as (first, second, third , fourth) You can do it ,right, right?

OpenStudy (anonymous):

of course :) thank you!

OpenStudy (loser66):

good